OpenStudy (anonymous):

solve using elimination method in an ordered pair:2x+3y=3 4x+6y=6

myininaya (myininaya):

multiply first equation by 2 and you will see they are the same line

jimthompson5910 (jim_thompson5910):

2x+3y=3 4x+6y=6 -2(2x+3y)=-2(3) 4x+6y=6 -4x-6y=-6 4x+6y=6 (-4x+4x)+(-6y+6y) = -6+6 0x+0y = 0 0 = 0 So there are an infinite number of solutions